OpenStudy (anonymous):

Please HELP!!! A student guesses the answers to all 10 questions on a true/false test. What is the probability that the student gets 6 questions correct? MULTIPLE CHOICE

OpenStudy (anonymous):

a. 207/531 b. 24/106 c. 105/512 d. 102/449

OpenStudy (amistre64):

what have you considered?

OpenStudy (amistre64):

id use the binomial distribution such that for a given number of outcomes, k, from a given set, n, the results are:\[\binom{n}{k}p^k(1-p)^{n-k}\]

OpenStudy (anonymous):

What is the n over k thing?

OpenStudy (amistre64):

its a binomial notation. it means n choose k

OpenStudy (amistre64):

in factorial notation its:\[\frac{n!}{k!(n-k)!}\]

OpenStudy (anonymous):

Ok and what is the p

OpenStudy (amistre64):

the probabilty of picking a correct solution.

OpenStudy (anonymous):

Ok I know what to do now.

OpenStudy (amistre64):

the math ....

OpenStudy (amistre64):

what is your probabilty of picking a correct answer? out of 2 choices?

OpenStudy (amistre64):

lol, i thought you asked 'what do i do now' :)

OpenStudy (anonymous):

the probability is 1/2 ... and ya haha

OpenStudy (amistre64):

\[\binom{10}{6}(1/2)^6(1/2)^{4}=\binom{10}{4}(.5)^{10}\] 10.9.8.7 ------- 4.3.2.1 10.9.7 ------ = 10*3*7 = 210 for the binomial part 3.1

OpenStudy (anonymous):

Is the answer 105/512 (c) @amistre64

OpenStudy (amistre64):

i get 105/512 as well

OpenStudy (anonymous):

ok thank you

OpenStudy (amistre64):

youre welcome